if modulation index is 1/2 and power of carrier wave is  2W.then what will be the total power in modulated wave? 

Grade:12

2 Answers

Beenu Mathew
7 years ago
Hi divya,
great question...there is a formula for that
Pm= Pc (1+m2/2)
Pm= 2 ( 1+ 1/8)
Pm= 9/4 W or 1.5W
 
Thank you,
BEST OF LUCK!!
 
JEHOYESJIRA
Shaurya Singh Chauhan
6 years ago
Pm= pc (1+ m^2 /2) WherePm- power of modulated wavePc- power of carrier wavem- modulation index = 2 (1+ 1/8) = 9/4Pm= 2.25 W